OpenStudy (anonymous):

simplify : x^2(y^2*z^3)^-2/y^3(xz^-3)^2

OpenStudy (anonymous):

um...... Wait

OpenStudy (diyadiya):

\[\huge \frac{x^2y^{-4}z^-{6}}{y^3x^2z^{-6}}\] Can you do now ?

OpenStudy (anonymous):

Actually, no . It's so confusing.

OpenStudy (anonymous):

How did you do the last result please ?

OpenStudy (diyadiya):

Remember :\[\huge x^m \times x^n = x^{m+n}\]\[\huge x^{-m}= \frac{1}{x^m}\]

OpenStudy (diyadiya):

\[\huge \frac{x^2(y^2*z^3)^{-2}}{y^3(xz^{-3})^2}=\frac{x^2 (y^2)^{-2}(z^3)^{-2}}{y^3(x)^2(z^{-3})^2}\]

OpenStudy (anonymous):

I have to say " thank you a lot Diyadiya "

OpenStudy (diyadiya):

You still have to simplify it You're Welcome (:

OpenStudy (anonymous):

atleast give her medals!!!

OpenStudy (anonymous):

Where are you ? I need your help again .

OpenStudy (anonymous):

Sorry " I forgot to click on Good answer "

OpenStudy (diyadiya):

No problem your other question has been answered by ishaan :)

OpenStudy (anonymous):

but it's not clear answer .

OpenStudy (diyadiya):

ok wait

OpenStudy (diyadiya):

sorry i gtg now ask in the chat for someone to help you out :)

OpenStudy (anonymous):

okay .